OpenStudy (anonymous):

Write an equivalent expression using radical notation? (a^8b^8)^1/9

OpenStudy (anonymous):

\[a^\frac{8}{9}b^\frac{8}{9}\]

OpenStudy (anonymous):

or if you prefer \[(ab)^\frac{8}{9}\]

OpenStudy (anonymous):

oops radical notation. try \[\sqrt[9]{a^8b^8}\]

OpenStudy (anonymous):

9√a^8b^8 with a little 9